# MHRC-Bench: A Multilingual Hardware Repository-Level Code Completion benchmark

### v1

@@ 4.3 Main Results
Table 4 shows the evaluation results of base models and fine-tuned models in different hardware languages. Prior to supervised fine-tuning (SFT), all models achieve nearly zero EM scores, reflecting the difficulty of repository-level hardware code completion for base models. After fine-tuning on our dataset, the performance of all models improves substantially, with fine-tuned variants significantly outperforming their base models. Overall, the post-trained Qwen2.5-Coder-14B consistently outperforms Qwen2.5-Coder-7B, which in turn achieves better performance than DS-Coder-6.7B and DS-Coder-5.7B, and substantially surpasses Qwen2.5-Coder-3B. We detail how SFT improves model performance in Appendix A.4.
@@
Table 5 reports the evaluation results of different retrieval methods. With the exception of GraphCoder, all retrieval approaches consistently improve model performance across multiple hardware languages. The reasons for GraphCoder’s inferior performance are discussed in Appendix A.5. Among the evaluated retrieval methods, RLcoder achieves the strongest overall performance. When comparing models using the same retrieval strategy, Gemini 2.5 Pro attains the best overall results, while GPT-5 exhibits better performance on specific metrics, including EM and BLEU for V/SV, as well as ES and BLEU for VHDL.
@@
Comparison between pretrained and fine-tuned models show that fine-tuned models consistently achieve higher EM, ES, and BLEU scores than larger, general-purpose models with substantially more parameters. For example, Qwen2.5-Coder-14B, after fine-tuning on our dataset, outperforms GPT-5, Grok-4, and DeepSeek V3.2 across all hardware languages, and surpasses Gemini 2.5 Pro in all cases except Chisel. These results demonstrate that our training dataset effectively enhances the hardware code completion capability of code LLMs, even with a smaller model scale.
4.4 Experimental Analysis
4.4.1 Analysis on CodeBLEU and Compilation rate.
CodeBLEU (Ren et al., 2020) evaluates code generation quality by considering not only surface-level n-gram overlap but also syntactic parsability and semantic consistency. In Table 6, we report CodeBLEU scores across four hardware description languages. In addition, Table 7 presents the compilation success rates for different languages. Together, these results show that fine-tuning substantially improves the performance of Qwen2.5-Coder-7B.
Table 7: Models’ compilation rate (%) across hardware description languages. The procedure for computing the compilation rate is described in Appendix A.7.
@@ 4.4 Experimental Analysis
4.4.2 Analysis on different training data sizes.
4.4.3 Analysis on different lines.
As shown in Figure 4(c), post-trained Qwen2.5-Coder-7B and GPT-5 can effectively complete tasks involving a small number of lines. However, as the number of lines to be completed increases, the scores of the generated code gradually decline. This indicates that completing multi-line code remains a challenge for code LLMs.
4.4.4 Analysis on the prompt length.
For a fair comparison across models, we set the default prompt length to 2048 tokens. Nevertheless, varying the prompt length leads to different performance trends for both fine-tuned and general-purpose models, as illustrated in Figure 5. We therefore evaluate prompt lengths of 512, 1024, 2048, and 4096 tokens. The results show that GPT-5 achieves consistently higher EM performance at 2048 and 4096 tokens than at shorter prompt lengths. In contrast, the fine-tuned Qwen2.5-Coder-7B performs better at 2048 tokens than at 4096 tokens. Based on these observations, we adopt 2048 tokens as the input prompt length, which provides strong and stable performance across models.
4.4.5 Analysis on code structural levels.
As mentioned in 3.5, we categorize our test cases into five types according to node depths in the concrete syntax tree. As shown in Figure 5(c), We present the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 on the test set across these different structural levels, and we observe that as the structural level decreases, the performance of LLM may either improve or degrade depending on the structure granularity, which is different from software shown in Liu et al. (2025b) where performance decreases consistently with increasing structural levels. For detailed performance across node types, please refer to Appendix A.8.
4.4.6 Analysis on hardware semantic levels.
Mentioned in section 3.5, we categorize the nodes within the concrete syntax tree into nine primary semantic levels based on their semantic characteristics, and we provide the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 for these various semantic levels across multilingual languages on the test set, as illustrated in Figure 6(a) and 6(b). Notably, we observe significant performance disparities across different semantic levels. Specifically, models show superior performance on “Declaration and Definition”, while exhibiting lower efficacy on “Control Flow Block”. This suggests that current code LLMs are proficient at completing tasks related to variable definitions and references, yet their capacity to handle complex codes are limited. Appendix A.9 details performance on different language’s semantic types.
@@ 4.5 Key Observations
Performance varies significantly across hardware languages. Model performance differs substantially across hardware description languages. SystemVerilog/Verilog and HLS consistently achieve the best performance across all evaluated LLMs, whereas VHDL exhibits the lowest. This gap highlights the uneven maturity of LLMs in handling different hardware design paradigms.
Post-training is critical for improving hardware code completion. Post-training substantially enhances LLM performance on hardware code completion tasks. After fine-tuning on MHRC-Bench-train, code models show significant improvements over their base models and are able to surpass several general-purpose models. This demonstrates the effectiveness of targeted post-training for hardware-oriented tasks.
@@
Differences between hardware and software code. (1) Structural sensitivity: For software code, deeper concrete syntax tree (CST) structures often correlate with improved LLM performance Liu et al. (2025b). In contrast, for hardware code, increasing structural complexity typically leads to performance degradation. (2) Prompt length sensitivity. Increasing prompt length generally benefits software code completion, whereas for hardware code it yields limited or inconsistent gains (from 2048 to 4096 tokens), reflecting the greater difficulty of reasoning about hardware semantics.
@@
Inappropriate retrieval can degrade performance. Retrieval-augmented methods are not always beneficial. We observe that retrieving irrelevant or noisy information can harm model performance, as exemplified by GraphCoder. This highlights the importance of retrieval quality and task-aware retrieval strategies.
Conclusion
@@
We present MHRC-Bench, the first multilingual, repository-level benchmark for hardware code completion, covering RTL, HLS, and generator-based hardware languages across both design and verification stages. By introducing fine-grained annotations at the code-structure and hardware-semantic levels, MHRC-Bench enables detailed analysis of LLM behavior in hardware-oriented settings. Extensive experiments show that post-training on MHRC-Bench-Train substantially improves hardware code completion performance, allowing relatively small open-weight models to outperform strong commercial baselines. Our findings highlight fundamental differences between hardware and software code completion and establish MHRC-Bench as a valuable resource for advancing LLMs in hardware design automation.
Limitations
@@
First, while MHRC-Bench covers multiple major hardware description paradigms, it does not yet include analog/mixed-signal or low-level physical design languages. Second, our evaluation focuses on textual similarity metrics, which may not fully capture functional correctness of generated hardware code. Third, MHRC-Bench does not cover analog/mixed-signal and physical design languages but mainly digital hardware description languages. Addressing these limitations is an important direction for future work.
Ethics Statement
MHRC-Bench is constructed entirely from publicly available GitHub repositories released under permissive open-source licenses (e.g., MIT, BSD, Apache). Repositories with restrictive or copyleft licenses are excluded, and duplicate repositories and vendor/dependency directories are removed. The dataset contains only source code and does not intentionally include personal or private information. MHRC-Bench covers multiple hardware languages and abstraction levels to reduce evaluation bias, though it reflects the distribution of open-source hardware projects and may not represent proprietary practices. The benchmark is intended solely for research on hardware code completion and model evaluation. Generated code should not be deployed without proper human review, simulation, and verification.
While the proposed system automates aspects of hardware design, we recognize that AI-generated design suggestions should always be validated by domain experts before deployment in safety-critical or commercial environments. We encourage responsible use and open scientific collaboration. Also, we use ChatGPT to polish our paper writing.
@@ Appendix A Appendix
A.1 Cross-file dependency
@@
A.2 Distribution across hardware sematic categories
@@
A.3 Evaluation Configurations
A.3.1 Prompts
@@
Task Prompt You are a precise code-completion model completing a repository blank completion task. Input: a repository concatenation where each file of the repo is wrapped in --- FILE: <repo_name>/<rel/path/to/file> (BEGIN) ---<file content...> --- FILE: <repo name>/<rel/path/to/file> (END) ---. Some lines have been replaced by the exact token <TARGET>. Your task: reconstruct only the replaced line, using the repository context. Note that for certain large repositories, functions of lesser importance were removed of their implementation, and only the signature/documentation were kept for the sake of simplicity.
OUTPUT: the reconstructed line(and nothing else).
@@
Figure 8: The prompt used for repository-level code completion (No RAG). If the context length exceed the token limit, the head of the left context and the tail of the right context are trimmed.
A.3.2 Fine-Tuning
For fine-tuning local LLMs (Qwen and Deepseek), the training objective was modeled as a "fill-in-the-middle" instruction task, where the model is prompted with a concatenated repository context containing file delimiters and a specific <TARGET> token representing the missing line. To manage context window constraints, inputs were dynamically truncated to a maximum of 2,048 tokens, employing a balanced budgeting strategy to prioritize the immediate left and right context surrounding the target. Fine-tuning was achieved using Low-Rank Adaptation (LoRA) applied to all linear attention and feed-forward layers, configured with a rank $r=16$, scaling factor $\alpha=32$, and a dropout of $0.05$. The model was optimized using AdamW with a peak learning rate of $2\times 10^{-4}$, following a cosine decay schedule with a 3% warm-up period all over a single epoch. Cross-entropy loss was computed exclusively on the generated ground-truth line, masking the instruction and context tokens. For reference, we plot the loss against the epoch number for Qwen2.5-Coder-7B, Qwen2.5-Coder-14B, and deepseek-coder-5.7bmqa-base in Figure 9.
A.3.3 Evaluation Metrics
@@
Prior to calculating the EM, ES and BLEU metrics, the generated code were stripped of comments, followed by trimming whitespace and dropping empty lines. The EM is set to $1$ if the cleaned generation results exactly matches the ground truth, otherwise it is set to 0. For ES, we evaluate the following formula:
@@
where editdistanceeditdistance is the Levenshtein edit distance between the generated code and the ground truth. $BLEU$ is calculated as the token-level BLEU between the cleaned generation results and the ground truth.
A.3.4 Commercial API setting
@@
A.4 Case Study of Pretrained-Model Responses
To illustrate the effect of fine-tuning on local models, we analyze the completions generated by Qwen2.5-Coder-7B before and after fine-tuning. Specifically, we list for each tested language common errors made by the model that were significantly improved by fine-tuning.
A.4.1 Verilog/SystemVerilog

### v2

@@ 4.3 Main Results
Table 4 reports the evaluation results of different retrieval methods with the default setting in their papers. With the exception of GraphCoder, all retrieval approaches consistently improve model performance across multiple hardware languages. The reasons for GraphCoder’s inferior performance are discussed in Appendix A.6. Among the evaluated retrieval methods, RLcoder achieves the strongest overall performance. When comparing models using the same retrieval strategy, Gemini 2.5 Pro attains the best overall results, while GPT-5 exhibits better performance on specific metrics, including EM for V/SV, as well as ES for VHDL. Table 5 summarizes the performance of base code models, fine-tuned models, and fine-tuned models augmented with the RLcoder retriever across different hardware description languages. Before supervised fine-tuning (SFT), all models achieve near-zero EM scores, underscoring the difficulty of repository-level hardware code completion for pretrained models. Fine-tuning yields substantial performance gains across all models, and retrieval augmentation further improves results. Overall, post-trained Qwen2.5-Coder-14B achieves the strongest performance, followed by Qwen2.5-Coder-7B, which outperforms DS-Coder-6.7B and DS-Coder-5.7B, while Qwen2.5-Coder-3B performs the worst. Further analysis of SFT effects is provided in Appendix A.5.
Comparison between pretrained and fine-tuned models shows that fine-tuned models consistently achieve higher EM and ES scores than larger, general-purpose models with substantially more parameters. For example, Qwen2.5-Coder-14B, after fine-tuning on our dataset, outperforms GPT-5, Grok-4, and DeepSeek V3.2 across all languages, and surpasses Gemini 2.5 Pro in all cases except Chisel.
4.4 Evaluation of Other Benchmarks and Models
We compare Verilog-specific models (Wu et al., 2025; Liu et al., 2024b) across a prior hardware repository benchmark and ours. To ensure a fair comparison in Table 6, all models use the same base model (DeepSeek-Coder-6.7B-base) and the preceding-context-only setting. We further remove 11 overlapping repositories between MHRC-train and RTL-repo from the MHRC training set to avoid data leakage. Despite the increased difficulty of MHRC-Eval due to multi-line completion tasks, MHRC-Deepseek achieves strong performance on this benchmark and surpasses RTLCoder-DeepSeek, which indicates that our training dataset has good generalization without overfitting.
4.5 Ablation Study
4.5.1 Analysis on context format.
As shown in Tables 4 and 5, the full-context outperforms the preceding-context-only in terms of EM and ES across all models and languages except VHDL. This suggests that bidirectional context is generally beneficial for code generation, while the rigid and highly structured nature of VHDL reduces the marginal benefit of right-context information. Hence, we focus on full-context in the following analysis.
(a) Ablation study of dataset size on performance for fine-tuning Qwen2.5-Coder-7B.
(b) Ablation study of model performance and the number of lines to be completed.
(c) Ablation study of prompt lengths: EM of GPT-5 and post-trained Qwen2.5-Coder-7B.
Figure 3: Ablation study of training data scale, line number, completion length on hardware code completion.
4.5.2 Analysis on different training data sizes.
Figure 4: Study of Model performance across different code structure depths.
4.5.3 Analysis on the prompt length.
For a fair comparison across models, we set the default prompt length to 2048 tokens. Nevertheless, varying the prompt length leads to different performance trends for both fine-tuned and general-purpose models, as illustrated in Figure 3(c). We therefore evaluate prompt lengths of 512, 1024, 2048, and 4096 tokens. The results show that GPT-5 achieves consistently higher EM performance at 2048 and 4096 tokens than at shorter prompt lengths. In contrast, the fine-tuned Qwen2.5-Coder-7B performs better at 2048 tokens than at 4096 tokens. Based on these observations, we adopt 2048 tokens as the input prompt length, which provides strong and stable performance across models.
4.5.4 Analysis on different lines.
As shown in Figure 3(b), post-trained Qwen2.5-Coder-7B and GPT-5 can effectively complete tasks involving a small number of lines. However, as the number of lines to be completed increases, the scores of the generated code gradually decline. This indicates that completing multi-line code remains a challenge for code LLMs.
Table 6: ES (%) and EM (%) of DeepSeek-based models evaluated on MHRC-Eval (V/SV) and RTL-repo test.
Benchmark Model ES EM
MHRC-Eval (V/SV) RTLCoder-DeepSeek(Liu et al., 2024b) 15.5 0.0
MHRC-Deepseek 63.7 28.5
RTLRepoCoder-DeepSeek(Wu et al., 2025) 13.0 6.3
RTL-repo Test RTLCoder-DeepSeek(Liu et al., 2024b) 48.1 16.2
MHRC-Deepseek 51.7 18.4
RTLRepoCoder-DeepSeek(Wu et al., 2025) 81.2 50.3
4.5.5 Analysis on code structural levels.
As mentioned in 3.5, we categorize our test cases into five types according to node depths in the concrete syntax tree. As shown in Figure 4, We present the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 on the test set across these different structural levels, and we observe that as the structural level decreases, the performance of LLM may either improve or degrade depending on the structure granularity, which is different from software shown in Liu et al. (2025b) where performance decreases consistently with increasing structural levels. For detailed performance across node types, please refer to Appendix A.9.
4.5.6 Analysis on hardware semantic levels.
Mentioned in section 3.5, we categorize the nodes within the concrete syntax tree into nine primary semantic levels based on their semantic characteristics, and we provide the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 for these various semantic levels across multilingual languages on the test set, as illustrated in Figure 5. Notably, we observe significant performance disparities across different semantic levels. Specifically, models show superior performance on “Declaration and Definition”, while exhibiting lower efficacy on “Control Flow Block” and ”Monitoring and Checking Logic”. This suggests that current code LLMs are proficient at completing tasks related to variable definitions and references, yet their capacity to handle complex code is limited. Appendix A.10 details performance on different language’s semantic types.
4.5.7 Analysis on CodeBLEU and Compilation Pass Rate.
CodeBLEU (Ren et al., 2020) evaluates code generation quality by considering not only surface-level n-gram overlap but also syntactic parsability and semantic consistency. In Table 7, we report CodeBLEU scores across four hardware description languages. In addition, Table 8 presents the compilation pass rates for different languages. Together, these results show that fine-tuning substantially improves the performance of Qwen2.5-Coder-7B.
Figure 5: EM performance of GPT-5 and post-trained Qwen2.5-Coder-7B across hardware semantic categories.
@@ 4.5 Ablation Study
Table 8: Models’ compilation pass@1 (%) across hardware description languages. The procedure for computing the compilation rate is described in Appendix A.8.
@@ 4.6 Key Observations
Performance varies significantly across hardware languages. Model performance differs substantially across hardware description languages. SystemVerilog/Verilog and HLS consistently achieve the best performance across all evaluated LLMs, whereas VHDL exhibits the lowest. Also, the full-context setting consistently outperforms the preceding-context-only setting in EM and ES across all models and languages except for VHDL. This gap highlights the uneven maturity of LLMs in handling different hardware design paradigms.
Post-training is critical for improving hardware code completion. Post-training markedly enhances the performance of LLMs on hardware code completion tasks. After fine-tuning on MHRC-Bench-train, code-specialized models achieve significant improvements over their base counterparts and even outperform several general-purpose models. This advantage stems from two key factors. First, general-purpose models tend to generate more code than required shown in Appendix A.11, whereas fine-tuned models produce outputs that better align with task-specific expectations. Second, hardware description languages are underrepresented in the pretraining data of general-purpose models. These results demonstrate the effectiveness of targeted post-training for hardware-oriented code completion tasks.
@@
Differences between hardware and software code. (1) Structural sensitivity: For software code, deeper CST depths often correlate with improved LLM performance (Liu et al., 2025b). However, for hardware code, increasing structural complexity may increase performance or decrease performance. (2) Prompt length sensitivity. Increasing prompt length generally benefits software code completion, whereas for hardware code it yields limited or inconsistent gains (from 2048 to 4096 tokens), reflecting the greater difficulty of reasoning about hardware semantics.
@@
Inappropriate retrieval can degrade performance. Retrieval-augmented methods are not always beneficial. We observe that retrieving irrelevant or noisy information can harm model performance such as GraphCoder. GraphCoder fundamentally struggles with HLS and HDL code because their core assumptions about code structure and semantics do not hold in hardware-oriented programming detailed in A.6. This highlights the importance of retrieval quality and specific retrieval strategies for hardware code.
5 Conclusion and Limitations
@@
We present MHRC-Bench, the first multilingual, repository-level benchmark for hardware code completion, covering RTL, HLS, and generator-based hardware languages across design and verification stages. Fine-grained structural and semantic annotations enable detailed analysis of LLM behavior in hardware-oriented settings. Experiments show that post-training on MHRC-Bench-Train substantially improves performance, allowing relatively small open-weight models to outperform strong general baselines. These results highlight fundamental differences between hardware and software code completion and establish MHRC-Bench as a valuable resource for hardware design automation.
@@
However, MHRC-Bench currently focuses on digital hardware description languages and does not cover analog/mixed-signal or physical design languages. In addition, our evaluation relies on textual similarity metrics, CodeBLEU and compilation pass rate, which may not fully reflect functional correctness. Addressing these limitations is left for future work.
Impact Statement
This work seeks to advance machine learning for electronic design automation (EDA) by enabling more effective hardware code completion with large language models. By providing a multilingual, repository-level benchmark and systematic analysis, this work supports research on improving hardware design efficiency, verification workflows, and tool automation. We do not foresee any direct negative societal impacts. The proposed benchmark and findings are intended for research use and may contribute to more accessible and efficient hardware development practices.
@@ Appendix A Appendix
A.1 Dataset Splits
To make sure no overlap repository and files between evaluation and training, we perform the following algorithm. After we choose one file from repository randomly, we conducted an exact-match leakage check via hashing shown in the following algorithm. Specifically, for each example we concatenate the left context, ground-truth span, and right context into a single string, normalize text (line endings and optional whitespace normalization), and compute a SHA-1 digest. We build a hash index over all training examples and flag any evaluation example whose digest appears in the training index; such overlaps are removed to prevent train–test leakage.
Algorithm 1 Leakage-Free Test Set Construction
1:Input: Training set $\mathcal{D}_{train}$, candidate test files $\mathcal{C}$
2:Output: Leakage-free test set $\mathcal{D}_{test}$
3: Build hash index $\mathcal{H}_{train}$ from $\mathcal{D}_{train}$ by hashing left-context ++ ground-truth ++ right-context
[…]
5:for each candidate file $c\in\mathcal{C}$ do
6:  Construct full-context string $s_{c}$
7:  Normalize $s_{c}$ (line endings, optional whitespace)
8:  Compute hash $h_{c}\leftarrow\mathrm{SHA1}(s_{c})$
9:  if $h_{c}\notin\mathcal{H}_{train}$ then
10:   $\mathcal{D}_{test}\leftarrow\mathcal{D}_{test}\cup\{c\}$
11:  else
12:   Discard $c$ (potential train–test leakage)
13:  end if
14:end for
15:return $\mathcal{D}_{test}$
A.2 Cross-file dependency
@@
A.3 Distribution across hardware sematic categories
@@
A.4 Evaluation Configurations
A.4.1 Prompt
@@
Task Prompt You are a precise code-completion model completing a repository blank completion task. Input: a repository concatenation where each file of the repo is wrapped in --- FILE: <repo_name>/<rel/path/to/file> (BEGIN) ---<file content...> --- FILE: <repo name>/<rel/path/to/file> (END) ---. Some lines have been replaced by the exact token <TARGET>. Your task: reconstruct only the replaced lines, using the repository context.
OUTPUT: the reconstructed lines(and nothing else).
@@
Figure 7: The full context prompt used for repository-level code completion (No RAG). If the context length exceed the token limit, the head of the left context and the tail of the right context are trimmed.
Task Prompt You are a precise code-completion model completing a repository blank completion task. Input: a repository concatenation where each file of the repo is wrapped in --- FILE: <repo_name>/<rel/path/to/file> (BEGIN) ---<file content...> --- FILE: <repo name>/<rel/path/to/file> (END) ---. Some lines have been replaced by the exact token <TARGET>. Your task: reconstruct only the replaced lines, using the repository context.
OUTPUT: the reconstructed lines (and nothing else). Target line count: target_lines. Output exactly target_lines line(s).
<Left Context>... ...</Left Context>
<TARGET>
Figure 8: The Preceding context prompt used for repository-level code completion (No RAG). If the context length exceed the token limit, the head of the left context and the tail of the right context are trimmed.
A.4.2 Fine-Tuning
For fine-tuning local LLMs (Qwen and Deepseek), the training objective was modeled as a completion instruction task, where the model is prompted with a concatenated repository context containing file delimiters and a specific <TARGET> token representing the missing line. To manage context window constraints, inputs were dynamically truncated to a maximum of 2,048 tokens, employing a balanced budgeting strategy to prioritize the immediate left and right context surrounding the target. Fine-tuning was achieved using Low-Rank Adaptation (LoRA) applied to all linear attention and feed-forward layers, configured with a rank $r=16$, scaling factor $\alpha=32$, and a dropout of $0.05$. The model was optimized using AdamW with a peak learning rate of $2\times 10^{-4}$, following a cosine decay schedule with a 3% warm-up period all over a single epoch. Cross-entropy loss was computed exclusively on the generated ground-truth line, masking the instruction and context tokens. For reference, we plot the loss against the epoch number for Qwen2.5-Coder-7B, Qwen2.5-Coder-14B, and deepseek-coder-5.7bmqa-base in Figure 9.
A.4.3 Evaluation Metrics
@@
Prior to calculating the EM and ESmetrics, the generated code were stripped of comments, followed by trimming whitespace and dropping empty lines. The EM is set to $1$ if the cleaned generation results exactly matches the ground truth, otherwise it is set to 0. For ES, we evaluate the following formula:
@@
where editdistanceeditdistance is the Levenshtein edit distance between the generated code and the ground truth.
A.4.4 Commercial API setting
@@
A.5 Case Study of Pretrained-Model Responses
To illustrate the effect of fine-tuning on local models, we analyze the completions generated by Qwen2.5-Coder-7B on fill-in-the-middle tasks before and after fine-tuning. Specifically, we list for each tested language common errors made by the model that were significantly improved by fine-tuning.
A.5.1 Verilog/SystemVerilog